# The Efficacy of Semantics-Preserving Transformations in Self-Supervised Learning for Medical Ultrasound

### v1

@@ IV-B Object Classification Task Evaluation
PE SimCLR BYOL $0.801$ $0.783$ $0.788$ $0.893$
SimCLR AugUS-O $0.810$ $0.800$ $0.785$ $0.865$
SimCLR AugUS-D $\mathbf{0.816}$ $\mathbf{0.808}$ $\mathbf{0.789}$ $\mathbf{0.897}$
ImageNet - $0.806$ $0.774$ $0.820$ $0.884$
We fine-tuned the pretrained models, allowing the backbone’s weights to be trainable in addition to the model head. Table VI gives the test set performance of the fine-tuned classifiers. We observed similar performance differences among the different augmentation pipelines, but note some additional findings. The model pretrained using AugUS-O on LUSData performed comparably against the other pipelines on AB, but exhibited extremely poor performance on the PE test set. Although it may appear that this model may have overfit to the training set, examination of training metrics revealed that training and validation metrics were close, with validation set AUC having been evaluated as $0.861$. Nonetheless, fine-tuned models that were pretrained with the BYOL and AugUS-O pipelines yielded strong performance on both tasks in LUSData.
TABLE VI: Test set metrics for classifiers fine-tuned for each of the AB and PE tasks.
Task Initial Weights Pipeline Accuracy Precision Recall AUC
AB SimCLR BYOL $\mathbf{0.941}$ $0.951$ $0.850$ $\mathbf{0.970}$
@@
Fine-tuned classifiers for the AB and PE tasks were also evaluated on the external portion of the LUSData test set. Most classifiers exhibited degraded performance on external data, compared to the local test set. Overall, the relative performance of the classifiers on the external test set was reflective of their performance on local test data. Unlike the local test evaluation, the network trained from scratch performed comparably to the SimCLR-pretrained models that utilized the AugUS-O and AugUS-D pipelines; however, the SimCLR-pretrained model that utilized the BYOL pipeline achieved the greatest test AUC by a margin of $0.029$. On the PE task, the classifier originating from the same pretrained model that utilized BYOL achieved the greatest AUC by a marging of $0.018$. Similar to the local test set, the pretrained models that incorporated the BYOL and AugUS-D pipeline achieved the greatest test AUC, while the pretrained model that utilized AugUS-O performed the worst.
TABLE VII: External test set metrics for classifiers fine-tuned for the AB and PE tasks.
Task Initial Weights Pipeline Accuracy Precision Recall AUC
AB SimCLR BYOL $\mathbf{0.751}$ $\mathbf{0.961}$ $0.556$ $\mathbf{0.883}$
@@
PE SimCLR BYOL $\mathbf{0.805}$ $0.838$ $0.861$ $\mathbf{0.898}$
@@
IV-C Diagnostic Classification Task Evaluation
@@
Linear classifiers were trained on the COVID training set and evaluated on the COVID test set. As shown in Table VIII, AugUS-O was observed to have the greatest test multiclass AUC, which was considered the primary metric of interest. Looking at the t-SNE visualizations in Fig. 5, AugUS-O corresponds to the only visualization where the representations for the COVID-19 Pneumonia and non-COVID-19 Pneumonia classes are clustered together.
Table IX provides test metrics for fine-tuned COVID classifiers. Again, AugUS-O exhibited the best performance. Moreover, fine-tuned models generally performed worse than the linear classifiers trained on feature vectors from SSL-pretrained models; they likely suffered from overfitting, as COVIDx-US is a smaller dataset.
Unlike the AB and PE tasks, models trained for COVID were pretrained using the LUSData dataset. We repeated the linear classification and fine-tuning experiments using models pretrained on the COVIDx-US training set. Tables VIII and IX respectively report the results of linear and fine-tuning evaluations. As was observed for backbones pretrained on LUSData, pretraining with the AugUS-O pipeline resulted in the greatest test set AUC.
The trends observed for the COVID task are different than those observed for the AB and PE tasks. Regardless of the augmentation pipeline, SimCLR-pretrained weights resulted in better performance than ImageNet-pretrained or random weight initialization. On object classification tasks, models pretrained using the BYOL and AugUS-D pipelines performed the best. However, on the diagnostic COVID task, AugUS-O performed best. Recall that AugUS-O was designed to retain semantic information, while both BYOL and AugUS-D contain the very impactful crop & resize (C&R) transform that can obscure large portions of the image. Object classification tasks require scale invariance, which is enforced by applying C&R during SSL pretraining. Diagnostic tasks, on the other hand, require global image context for interpreters to make a decision, which is preserved best by the AugUS-O pipeline.
TABLE VIII: Test set performance for linear classifiers trained for the COVID task. Binary metrics are averages across classes.
Pretraining Dataset Initial Weights Pipeline Accuracy Precision Recall AUC
LUSData SimCLR BYOL $0.454$ $0.371$ $0.413$ $0.784$
@@
COVIDx-US SimCLR BYOL $0.498$ $\mathbf{0.582}$ $0.501$ $0.781$
@@
TABLE IX: Test set performance for fine-tuned classifiers trained for the COVID task. Binary metrics are averages across classes.
Pretraining Dataset Initial Weights Pipeline Accuracy Precision Recall AUC
LUSData SimCLR BYOL $0.381$ $0.259$ $0.365$ $0.753$
@@
COVIDx-US SimCLR BYOL $0.450$ $\mathbf{0.540}$ $0.464$ $0.770$
@@
IV-D Object Detection Task Evaluation
Recall that the PL task is an object detection problem geared toward localizing the pleural line artifact. We evaluated the pretrained models on PL to explore whether the trends observed for object-centric lung US classification tasks would hold for an object detection task, where locality understanding is explicit. We considered two evaluation settings: one in which the pretrained backbones weights were held constant, and another in which the backbone’s weights were trainable. Table X reports the average precision at a $50\%$ intersection over union threshold (AP@50) evaluated on the LUSData test set. When the backbone weights were frozen, SimCLR pretraining with BYOL and AugUS-D resulted in the greatest test AP@50. These trends reflected the results observed for AB and PE classification, which both require object recognition. However, when training end-to-end, models trained from scratch matched the performance of pretrained models.
TABLE X: LUSData local test set AP@50 for the PL task observed for SSD models whose backbones were pretrained using different data augmentation pipelines.
@@ IV-D Object Detection Task Evaluation
Frozen SimCLR BYOL $\mathbf{0.136}$
SimCLR AugUS-O $0.047$
SimCLR AugUS-D $0.108$
Random - $0.012$
ImageNet - $0.043$
Trainable SimCLR BYOL $0.136$
SimCLR AugUS-O $0.141$
SimCLR AugUS-D $\mathbf{0.143}$
Random - $0.142$
@@
IV-E Label Efficiency Assessment
Experiments were conducted to test the robustness of pretrained models in settings where few labeled samples are available. The experiment was conducted only for the AB and PE classification tasks because there were enough unique videos and patients in the training set to create several disjoint subsets. Backbones were fine-tuned on $20$ subsets of approximately $5\%$ of the LUSData training set, split by patient, yielding $20$ performance estimates for low-label settings. Splitting was conducted at the patient level to heighten the difficulty of the task and to limit dependence between subsets. Baseline estimates without SSL pretraining were obtained via initialization with random weights and with ImageNet-pretrained weights, resulting in five different performance conditions. Friedman’s Test indicated that there were significant differences among the median test AUCs across conditions, for both the AB and PE tasks. Post-hoc Wilcoxon Sign-Rank Tests were then conducted for each pair of conditions, using the Bonferroni correction with a family-wise error rate of $\alpha=0.05$. The median test AUCs of SimCLR-pretrained models were significantly greater than those initialized with random or ImageNet-pretrained weights for both the AB and PE tasks. All medians were significantly different for AB, except for the SimCLR-pretrained models using the BYOL and AugUS-D pipelines, which achieved the greatest performance. Notably, these pipelines both utilize the crop and resize transformation. No significant differences were observed between any of the SimCLR-pretrained models for PE. Appendix H provides the test statistics for the above comparisons.
(a) AB
(b) PE
@@ IV-E Label Efficiency Assessment
IV-F Impact of Semantic-Preserving Preprocessing
As outlined in Section III-B, all US images were cropped to the smallest rectangle enclosing the FOV because the areas outside the FOV are bereft of information. Since pipelines containing the crop and resize transformation (C&R) would be more likely to result in positive pairs that do not cover the FOV, it was hypothesized that cropping to the FOV as a preprocessing step would result in stronger pretrained backbones. To investigate the effect of this semantics-preserving preprocessing step, we pretrained backbones on LUSData using each data augmentation pipeline and evaluated them on the AB, AB, and COVID tasks. Table XI compares the performance of each backbone with and without the preprocessing step. Performance on the AB task did not change. However, test AUC on both the PE and COVID tasks was consistently lower when the semantics-preserving preprocessing was not applied. Note that greatly less labelled data is available for PE and COVID than for AB. Based on these experiments, FOV cropping is a valuable semantics-preserving preprocessing step for multiple LUS classification tasks.
TABLE XI: Test set AUC for SimCLR-pretrained models with and without semantics-preserving preprocessing. Results are reported for linear classifiers and fine-tuned models.
@@ IV-F Impact of Semantic-Preserving Preprocessing
IV-G Impact of the Cropping in Object Classification Tasks
The leave-one-out analysis for transformations exhibited the striking finding that crop and resize (C&R) was the most effective transformation in the BYOL pipeline for the two object classification tasks: AB and PE. Moreover, both pipelines containing C&R resulted in the greatest downstream test performance on AB and PE. Ordinarily, crops are taken at random locations in an image, with areas between $8\%$ and $100\%$ of the original image’s area. Aggressive crops can create situations in which positive pairs do not contain the same objects of interest. For example, Fig. 7 shows how C&R could produce a positive pair where the views would possess different labels. Despite this, the results indicated that pipelines containing C&R led to improved performance for the object-centric AB and PE tasks. The exceptional influence of C&R warranted further investigation into its hyperparameters.
Figure 7: Examples of how the random crop and resize transformation (B00) can reduce semantic information. Original images are on the left, and two random crops of the image are on the right. Top: The original image contains a B-line (purple), which is visible in View 2, but not in View 1. Bottom: The original image contains a pleural effusion (green), which is visible in View 1, but is largely obscured in View 2.
We investigated the impact of the minimum crop area, $c$, as a hyperparameter. Models were pretrained with the AugUS-D pipeline, using values for $c$ in the range $[0.05,0.9]$. Linear evaluation was conducted for the AB and PE tasks. As shown in Fig. 8, smaller values of $c$ yielded better performance, peaking at $c\approx 0.1$. The default assignment of $c=0.08$ was already a reasonable choice for these two tasks. Additional experiments elucidating the effects of C&R hyperparameters can be found in Appendix I.
Another concern with C&R is that it could result in crops covering the black background on images with a convex FOV. Despite the semantics-preserving preprocessing (described in Section 2), the top left and right corners of such images provide no information. To characterize the robustness of pretraining under these circumstances, we repeated the experiments sweeping over $c\in[0.05,0.9]$, but first applied the probe type change transformation (i.e., U00) to every convex FOV. Thus, all inputs to the model were linear FOVs devoid of non-semantic background. A by-product of this transformation is that the near fields of convex images are horizontally stretched. As seen in Fig. 8, this change resulted in a slight decrease in performance for both the AB and PE tasks. Evidently, the detriment of spatial distortion outweighed the benefit of guaranteeing that crops were positioned over semantic regions.
(a) AB
(b) PE
@@ IV-G Impact of the Cropping in Object Classification Tasks
V Conclusion
This study proposed data augmentation and preprocessing strategies for self-supervised learning in ultrasound. A commonly employed baseline pipeline (BYOL) was compared to a handcrafted semantics-preserving pipeline (AugUS-O) and a hybrid pipeline (AugUS-D) composed from the first two. Evaluation on lung US interpretation tasks revealed a dichotomy between the utility of the pipelines. Pipelines featuring the cropping transformation (BYOL and AugUS-D) were most useful for object classification and detection tasks in LUS. On the other hand, AugUS-O – designed to preserve semantics in LUS – resulted in the greatest performance on a diagnostic task that required global context. Additionally, US field of view cropping was found to be a beneficial preprocessing step for multiple lung US classification tasks, regardless of the data augmentation strategy. Based on the results, practitioners should use semantics-preserving preprocessing during pretraining. Regarding data augmentation, semantics-preserving transformations should be considered for tasks requiring holistic interpretation of images, cropping-based transformations should be leveraged for object-centric downstream tasks.

### v2

@@ 4.2 Object Classification Task Evaluation
PE SimCLR StandardAug $0.782$ $0.769$ $0.787$ $0.881$
SimCLR AugUS-O $0.795$ $0.796$ $0.776$ $0.853$
SimCLR AugUS-D $\mathbf{0.800}$ $\mathbf{0.798}$ $0.782$ $\mathbf{0.886}$
ImageNet - $0.779$ $0.756$ $\mathbf{0.804}$ $0.864$
FT AB SimCLR StandardAug $\mathbf{0.941}$ $0.951$ $0.850$ $\mathbf{0.970}$
@@
We fine-tuned the pretrained models, allowing the backbone’s weights to be trainable in addition to the model head. Table 5 gives the test set performance of the fine-tuned classifiers. We observed similar performance differences among the different augmentation pipelines, but note some additional findings. The model pretrained using AugUS-O on LUSData performed comparably against the other pipelines on AB but exhibited extremely poor performance on the PE test set. Although it may appear that this model may have overfit to the training set, examination of training metrics revealed that training and validation metrics were close, with validation set AUC having been evaluated as $0.861$. Nonetheless, fine-tuned models that were pretrained with the StandardAug and AugUS-O pipelines yielded strong performance on both tasks in LUSData.
Fine-tuned classifiers for the AB and PE tasks were also evaluated on the external portion of the LUSData test set. Most classifiers exhibited degraded performance on external data, compared to the local test set. Overall, the relative performance of the classifiers on the external test set was reflective of their performance on local test data. Unlike the local test evaluation, the network trained from scratch performed comparably to the SimCLR-pretrained models that utilized the AugUS-O and AugUS-D pipelines; however, the SimCLR-pretrained model that utilized the StandardAug pipeline achieved the greatest AB test AUC by a margin of $0.029$. On the PE task, the classifier originating from the same pretrained model that utilized StandardAug achieved the greatest AUC by a margin of $0.015$. Similar to the local test set, the pretrained models that incorporated the StandardAug and AugUS-D pipeline achieved the greatest test AUC, while the pretrained model that utilized AugUS-O performed the worst.
Although MobileNetV3Small was the backbone architecture used for these experiments, we repeated the above evaluations using the more commonly employed ResNet18 architecture [40]. Similar trends were observed regarding the greater test performance attained by models pretrained with cropping-based pipelines. However, the fine-tuned models greatly overfit, likely due to ResNet18’s much greater capacity. The ResNet18 models that achieved the greatest test performance were the linear classifiers trained on frozen backbones. Notably, the trend persisted when evaluating on external test data. Detailed results for ResNet18 can be found in H.
Table 6: External test set metrics for linear classification (LC) and fine-tuning (FT) experiments with the AB and PE tasks. Binary metrics are averages across classes. The best observed metrics in each experimental setting are in boldface.
Train Setting Task Initial Weights Pipeline Accuracy Precision Recall AUC
LC AB SimCLR StandardAug $\mathbf{0.749}$ $\mathbf{0.956}$ $\mathbf{0.555}$ $\mathbf{0.868}$
SimCLR AugUS-O $0.689$ $0.927$ $0.453$ $0.810$
SimCLR AugUS-D $0.726$ $0.922$ $0.531$ $0.859$
ImageNet - $0.643$ $0.872$ $0.389$ $0.770$
PE SimCLR StandardAug $0.794$ $0.835$ $\mathbf{0.843}$ $0.880$
SimCLR AugUS-O $0.784$ $\mathbf{0.916}$ $0.728$ $0.870$
SimCLR AugUS-D $\mathbf{0.806}$ $0.877$ $0.809$ $\mathbf{0.887}$
ImageNet - $0.758$ $0.836$ $0.771$ $0.840$
FT AB SimCLR StandardAug $\mathbf{0.751}$ $\mathbf{0.961}$ $0.556$ $\mathbf{0.883}$
@@
PE SimCLR StandardAug $\mathbf{0.805}$ $0.838$ $0.861$ $\mathbf{0.898}$
@@
4.3 Diagnostic Classification Task Evaluation
@@
Linear classifiers were trained on the COVID training set and evaluated on the COVID test set. As shown in Table 7, AugUS-O was observed to have the greatest test multiclass AUC, which was considered the primary metric of interest. Looking at the t-SNE visualizations in Fig. 5, AugUS-O corresponds to the only visualization where the representations for the COVID-19 Pneumonia and non-COVID-19 Pneumonia classes are clustered together.
Table 7 provides test metrics for fine-tuned COVID classifiers. Again, AugUS-O exhibited the best performance. Moreover, fine-tuned models generally performed worse than the linear classifiers trained on feature vectors from SSL-pretrained models; they likely suffered from overfitting, as COVIDx-US is a smaller dataset.
Unlike the AB and PE tasks, models trained for COVID were pretrained using the LUSData dataset. We repeated the linear classification and fine-tuning experiments using models pretrained on the COVIDx-US training set. Table 7 reports the results of linear and fine-tuning evaluations. As was observed for backbones pretrained on LUSData, pretraining with the AugUS-O pipeline resulted in the greatest test set AUC.
The trends observed for the COVID task are different than those observed for the AB and PE tasks. Regardless of the augmentation pipeline, SimCLR-pretrained weights resulted in better performance than ImageNet-pretrained or random weight initialization. On object classification tasks, models pretrained using the StandardAug and AugUS-D pipelines performed the best. However, on the diagnostic COVID task, AugUS-O performed best. Recall that AugUS-O was designed to retain semantic information, while both StandardAug and AugUS-D contain the very impactful crop & resize (C&R) transform that can obscure large portions of the image. Object classification tasks require scale invariance, which is enforced by applying C&R during SSL pretraining. Diagnostic tasks, on the other hand, require global image context for interpreters to make a decision, which is preserved best by the AugUS-O pipeline.
Table 7: Test set performance for linear classification (LC) and fine-tuning (FT) experiments with the COVID task. Binary metrics are averages across classes. The best observed metrics in each experimental setting are in boldface.
Train Setting Pretraining Dataset Initial Weights Pipeline Accuracy Precision Recall AUC
LC LUSData SimCLR StandardAug $0.454$ $0.371$ $0.413$ $0.784$
@@
COVIDx-US SimCLR StandardAug $0.498$ $\mathbf{0.582}$ $0.501$ $0.781$
@@
FT LUSData SimCLR StandardAug $0.381$ $0.259$ $0.365$ $0.753$
@@
COVIDx-US SimCLR StandardAug $0.450$ $\mathbf{0.540}$ $0.464$ $0.770$
@@
4.4 Object Detection Task Evaluation
Recall that the PL task is an object detection problem geared toward localizing the pleural line. We evaluated the pretrained models on PL to explore whether the trends observed for object-centric lung US classification tasks would hold for an object detection task, where locality understanding is explicit. We considered two evaluation settings: one in which the pretrained backbones weights were held constant, and another in which the backbone’s weights were trainable. Table 8 reports the average precision at a $50\%$ intersection over union threshold (AP@50) evaluated on the LUSData test set. When the backbone weights were frozen, SimCLR pretraining with AugUS-O resulted in the greatest test AP@50. These trends differed from the results observed for AB and PE classification, which both require object recognition. We speculate that aggressive cropping during the pretraining phase likely produced positive pairs where one image contained a pleural line while the other did not, which we believe would make it difficult to learn representations for pleural line objects when pretraining with the StandardAug or AugUS-D pipelines. The performance of the frozen pretrained backbones was strong overall, considering the low capacity of the backbone and the small, narrow shape of pleural line objects. When fine-tuning end-to-end, SimCLR pretraining with AugUS-D resulted in the greatest test AP@50.
Table 8: LUSData local test set AP@50 for the PL task observed for SSD models whose backbones were pretrained using different data augmentation pipelines.
@@ 4.4 Object Detection Task Evaluation
Frozen SimCLR StandardAug $0.228$
SimCLR AugUS-O $\mathbf{0.255}$
SimCLR AugUS-D $0.194$
Random - $0.041$
@@
Trainable SimCLR StandardAug $0.316$
SimCLR AugUS-O $0.332$
SimCLR AugUS-D $\mathbf{0.351}$
Random - $0.308$
ImageNet - $0.310$
4.5 Label Efficiency Assessment
Experiments were conducted to test the robustness of pretrained models in settings where few labeled samples are available. The experiment was conducted only for the AB and PE classification tasks because there were enough unique videos and patients in the training set to create several disjoint subsets. Backbones were fine-tuned on $20$ subsets of approximately $5\%$ of the LUSData training set, split by patient, yielding $20$ performance estimates for low-label settings. Splitting was conducted at the patient level to heighten the difficulty of the task and to limit dependence between subsets. Baseline estimates without SSL pretraining were obtained via initialization with random weights and with ImageNet-pretrained weights, resulting in five different performance conditions. Figure 6 displays boxplots for test AUC distributions under each condition. Friedman’s Test indicated that there were significant differences among the median test AUC across conditions, for both the AB and PE tasks. Post-hoc Wilcoxon Sign-Rank Tests were then conducted for each pair of conditions, using the Bonferroni correction with a family-wise error rate of $\alpha=0.05$. The median test AUCs of SimCLR-pretrained models were significantly greater than those initialized with random or ImageNet-pretrained weights for both the AB and PE tasks. All medians were significantly different for AB, except for the SimCLR-pretrained models using the StandardAug and AugUS-D pipelines, which achieved the greatest performance. Notably, these pipelines both utilize the crop and resize transformation. No significant differences were observed between any of the SimCLR-pretrained models for PE. I provides the test statistics for the above comparisons.
(a) (a) AB classification task
(b) (b) PE classification task
@@ 4.5 Label Efficiency Assessment
4.6 Impact of Semantics-Preserving Preprocessing
As outlined in Section 3.2, all US images were cropped to the smallest rectangle enclosing the FOV because the areas outside the FOV are bereft of information. Since pipelines containing the crop and resize transformation (C&R) would be more likely to result in positive pairs that do not cover the FOV, it was hypothesized that cropping to the FOV as a preprocessing step would result in stronger pretrained backbones. To investigate the effect of this semantics-preserving preprocessing step, we pretrained backbones on LUSData using each data augmentation pipeline and evaluated them on the AB, AB, and COVID tasks. Table 9 compares the performance of each backbone with and without the preprocessing step. Performance on the AB task did not change. However, test AUC on both the PE and COVID tasks was consistently lower when the semantics-preserving preprocessing was not applied. Note that greatly less labelled data is available for PE and COVID than for AB. Based on these experiments, FOV cropping is a valuable semantics-preserving preprocessing step for multiple LUS classification tasks.
Table 9: Test set AUC for SimCLR-pretrained models with (✓) and without (✗) semantics-preserving preprocessing. Results are reported for linear classifiers and fine-tuned models.
@@ 4.6 Impact of Semantics-Preserving Preprocessing
4.7 Impact of the Cropping in Object Classification Tasks
The leave-one-out analysis for transformations exhibited the striking finding that crop and resize (C&R) was the most effective transformation in the StandardAug pipeline for the two object classification tasks: AB and PE. Moreover, both pipelines containing C&R resulted in the greatest downstream test performance on AB and PE. Ordinarily, crops are taken at random locations in an image, with areas between $8\%$ and $100\%$ of the original image’s area. Aggressive crops can create situations in which positive pairs do not contain the same objects of interest. Fig. 7 shows how C&R could produce positive pairs with different semantic content. Despite this, the results indicated that pipelines containing C&R led to improved performance for the object-centric AB and PE tasks. The exceptional influence of C&R warranted further investigation into its hyperparameters.
Figure 7: Examples of how the random crop and resize transformation (B00) can reduce semantic information. Original images are on the left, and two random crops of the image are on the right. Top: The original image contains a B-line (purple), which is visible in View 2 but not in View 1. The original image also contains instances of the pleural line (yellow) which are visible in View 1 but not in View 2. Bottom: The original image contains a pleural effusion (green), which is visible in View 1 but largely obscured in View 2.
We investigated the impact of the minimum crop area, $c$, as a hyperparameter. Models were pretrained with the AugUS-D pipeline, using values for $c$ in the range $[0.05,0.9]$. Linear evaluation was conducted for the AB and PE tasks. As shown in Fig. 8, smaller values of $c$ yielded better performance, peaking at $c\approx 0.1$. The default assignment of $c=0.08$ was already a reasonable choice for these two tasks. Additional experiments elucidating the effects of C&R hyperparameters can be found in J.
Another concern with C&R is that it could result in crops covering the black background on images with a convex FOV. Despite the semantics-preserving preprocessing (described in Section 2), the top left and right corners of such images provide no information. To characterize the robustness of pretraining under these circumstances, we repeated the experiments sweeping over $c\in[0.05,0.9]$ but first applied the probe type change transformation (i.e., U00) to every convex FOV. Thus, all inputs to the model were linear FOVs devoid of non-semantic background. A by-product of this transformation is that the near fields of convex images are horizontally stretched. As seen in Fig. 8, this change resulted in a slight decrease in performance for both the AB and PE tasks. Evidently, the detriment of spatial distortion outweighed the benefit of guaranteeing that crops were positioned over semantic regions.
(a) (a) AB classification task
(b) (b) PE classification task
@@ 4.7 Impact of the Cropping in Object Classification Tasks
5 Conclusion
This study proposed data augmentation and preprocessing strategies for self-supervised learning in ultrasound. A commonly employed baseline pipeline (StandardAug) was compared to a handcrafted semantics-preserving pipeline (AugUS-O) and a hybrid pipeline (AugUS-D) composed from the first two. Evaluation on lung US interpretation tasks revealed a dichotomy between the utility of the pipelines. Pipelines featuring the cropping transformation (StandardAug and AugUS-D) were most useful for object classification and detection tasks in LUS. On the other hand, AugUS-O – designed to preserve semantics in LUS – resulted in the greatest performance on a diagnostic task that required global context. Additionally, US field of view cropping was found to be a beneficial preprocessing step for multiple lung US classification tasks, regardless of the data augmentation strategy. Based on the results, developers should use semantics-preserving preprocessing during pretraining. Regarding data augmentation, semantics-preserving transformations should be considered for tasks requiring holistic interpretation of images, cropping-based transformations should be leveraged for object-centric downstream tasks.